OpenStudy (anonymous):

Find the slope and y-intercept of the graph of the equation. Show your work. 2y - 6 = 3x

OpenStudy (daniel.ohearn1):

First isolate y

OpenStudy (daniel.ohearn1):

Slope is m in the equation y=mx+b

OpenStudy (daniel.ohearn1):

y intercept will be where x=0

OpenStudy (anonymous):

First,convert this eqn\[2y-6=3x\]into this form\[y=mx+c\]where\[y=vertical~axis\]\[x=horizontal~axis\]\[m=gradient\]\[c=constant~or~y-intercept\]

OpenStudy (anonymous):

Hint: If u want to find y-intercept,\[x=0\] If u want to find x-intercept,\[y=0\]

OpenStudy (anonymous):

\[First,add~6~on~both~sides\]\[Then,~divide~both~sides~by~2\]

OpenStudy (anonymous):

\[2y-6+6=3x+6\]\[2y=3x+6\]

OpenStudy (anonymous):

\[\frac{ 2y }{ 2 }=\frac{ 3x+6 }{ 2 }\]\[y=\frac{ 3 }{ 2 }x+3\]

OpenStudy (anonymous):

Now,try comparing this eqn\[y=\frac{ 3 }{ 2 }x+3\]with this eqn\[y=mx+c\]

OpenStudy (anonymous):

where \[y=vertical~axis\]\[x=horizontal~axis\]\[m=gradient/slope\]\[c=constant/y-intercept\]

OpenStudy (anonymous):

@llama459

OpenStudy (anonymous):

Refer to the attached plot from Mathematica.